OpenStudy (anonymous):

Mary takes three hours to mow the yard. It takes Sally five hours. How long would it take with them mowing it together? Can someone help me work my way to the answer please? thanks

OpenStudy (anonymous):

A comprehensive approach would be to determine the rate at which Mary mows the yard and the rate at which Sally mows the yard. Then add up the two rates. Based on the information, at what rate does Mary mow a yard ?

OpenStudy (anonymous):

Think of both of them in terms of how much they get done in an hour. Mary mows 1/3rd of the lawn in an hour Sally mows 1/5th of the lawn in an hour How much of the lawn can they mow together in an hour? How many times must they do that to mow the entire lawn?

OpenStudy (anonymous):

so i would ad the 1/3 and 1/5 together??

OpenStudy (anonymous):

@blondie16

OpenStudy (anonymous):

If you add up the two rates you get a combined rate in "yards per hour". Swap the fraction to get "hours per yard".

OpenStudy (anonymous):

wait so i would do 3/1 +5/1 ??

OpenStudy (anonymous):

What did you get for the two rates? @mathdude69 told you what they were, so you just have to add them up :)

OpenStudy (anonymous):

so 8/2?

OpenStudy (anonymous):

Let's see . . . Mary's rate is 1/3 yard per hour Sally's rate is 1/5 yard per hour So their combined rate is: 1/3 + 1/5 = 8/15 The combined rate they both mow a lawn is 8/15 yard per hour Or, equivalently, 15/8 hours per yard = 1.875 hours per yard.

OpenStudy (anonymous):

wait how doe we get the 15?

OpenStudy (anonymous):

Because 1/3 + 1/5 is 8/15. You cannot add fractions unless they have a common denominator. \[\frac{ 1 }{ 3 } = \frac{ 5 }{ 15 }\] \[\frac{ 1 }{ 5 } = \frac{ 3 }{ 15 }\] \[\frac{ 1 }{ 3} + \frac{ 1 }{ 5 } = \frac{ 5 }{ 15 } + \frac{ 3 }{ 15 } = \frac{ 8 }{ 15 }\]

OpenStudy (anonymous):

oh okay thank you